OpenStudy (anonymous):

A number minus 20 times its reciprocal is 1. Find the number. If the problem has two solutions

OpenStudy (anonymous):

1

jhonyy9 (jhonyy9):

x - (20/x) =1 x2 - 20 =x x2 -x -20 =0 x_1,_2 =(1+/- sqrt(1+80))/2 =(1+/- 9 )/2 = -4 and 5 - prove -4 - (20/(-4))=1 -4-(-5)=1 -4+5=1 1=1 5 - (20/5)=1 5 - 4=1 1=1

jhonyy9 (jhonyy9):

- hope so much that will be right